OpenStudy (anonymous):

HELPPPPPPPP I'M DIEING!!! TO FINISH THIS, PLUS I GIVE MEDAL AND FAN PEEPZ!!!!!!!!!!!!!!!!!!!!!!..

Nnesha (nnesha):

i prefer chocolates

OpenStudy (anonymous):

Chocolate?

Nnesha (nnesha):

do yo know the formula ??

OpenStudy (anonymous):

Yes i know the formula

OpenStudy (abhisar):

Where is the sphere? What is its radius?

Nnesha (nnesha):

^^^^^^^^^^^^^^^^^^^^^^^

Nnesha (nnesha):

diagram ??

OpenStudy (anonymous):

V= 4/3 pi r^3

OpenStudy (anonymous):

umm hold on sorry

OpenStudy (anonymous):

Wait wrong question .-.

OpenStudy (abhisar):

:/

Nnesha (nnesha):

\(\rm\color{green}{o^\wedge\_^\wedge o}\)

OpenStudy (abhisar):

Do you know the formula?

OpenStudy (anonymous):

But V = 4/3 pir^3 is the formula

OpenStudy (anonymous):

Yes

OpenStudy (abhisar):

Formula is \(\boxed{\sf \huge \pi r^2\frac{h}{3}}\)

OpenStudy (anonymous):

What is the exact volume of the sphere? A. 2481.156 cm3 B. 7263.39 cm3 C. 7790.83 cm3 D. 22807.05 cm3 http://static.k12.com/calms_media/media/1566000_1566500/1566229/1/a08c85ca10d913bd3330f27923d726f23252e1f5/MS_IMC-141202-100703.jpg xD okay for sure thas the real one o.o

OpenStudy (anonymous):

:))

Nnesha (nnesha):

\(\Huge\ddot{\Large\frown}\)\(\rm\color{green}{o^\wedge\_^\wedge o}\) r you sure ??

OpenStudy (anonymous):

Positive

OpenStudy (abhisar):

ok, the last one is easy. The radius is given 12.3 cm. Simply plug in this value in the formula for volume of sphere and find the volume.

OpenStudy (anonymous):

This one don't have a height it just has a radius

OpenStudy (anonymous):

which is 12.3

OpenStudy (abhisar):

Yes, 12.3 is radius.

OpenStudy (anonymous):

I did 12.3 times 12.3 times 12.3 and got 1860

OpenStudy (abhisar):

\(\sf \huge \frac{4}{3} \times \pi \times (12.3)^3\)

OpenStudy (anonymous):

Well, I'm gonnna get 1860 when i times (12.3)^3 so its gonna look like....

OpenStudy (anonymous):

4/3 times pi times 1860... Correct?

OpenStudy (abhisar):

Yes.

OpenStudy (anonymous):

Okay i don't understand the fraction part :/ thats what i needed help on

OpenStudy (anonymous):

Wants i times my pi = 3.14 by 1860 i'm lost with the fraciton o.o

OpenStudy (anonymous):

fraction*

OpenStudy (abhisar):

First multiply 4*3.14*1860 and then divide the result with 3

OpenStudy (anonymous):

3.14 * 1860 = 5840

OpenStudy (anonymous):

@Abhisar

OpenStudy (abhisar):

Yes,

OpenStudy (anonymous):

Then i'm stuck with 4/3 times 5840

OpenStudy (anonymous):

How do i solve the fraction with that?

OpenStudy (anonymous):

@Abhisar

OpenStudy (abhisar):

Multiply 5840 with 4 and then divide the product with 3

OpenStudy (anonymous):

23360

OpenStudy (anonymous):

@Abhisar <--- ya ditching me :o

OpenStudy (abhisar):

:O ;-;

OpenStudy (abhisar):

I have explained you the process, now it's upto you to perform the calculations correctly.

Nnesha (nnesha):

yes that is right now* like kumar said divide by 3

OpenStudy (anonymous):

7786

OpenStudy (anonymous):

Aww the babies are so cute

OpenStudy (anonymous):

Anyone...

OpenStudy (anonymous):

There....

OpenStudy (anonymous):

-.- Thanx for the help

OpenStudy (anonymous):

23360 divide by 3 right?

Nnesha (nnesha):

wait nope there is righ answer

OpenStudy (anonymous):

I said 23360 and you said it was right

Nnesha (nnesha):

what did you use for pi ?? did yo just plug that in to the calculator or did you sue 3.14

OpenStudy (anonymous):

Used calculator o.o

Nnesha (nnesha):

yes that's why use 3.14

Nnesha (nnesha):

(12.3)^3 = ???

OpenStudy (anonymous):

1860

Nnesha (nnesha):

\[\large\rm \frac{ 4 }{ 3 }(3.14)(12.3)^3\]

Nnesha (nnesha):

yes now multiply that by 3.14 and then by 4

OpenStudy (anonymous):

5840

Nnesha (nnesha):

divide by 3

Nnesha (nnesha):

wait u did mistake somewhre

OpenStudy (anonymous):

Where

Nnesha (nnesha):

12.3^3 = is not just 1860

Nnesha (nnesha):

12.3^3 = 1860.867

Nnesha (nnesha):

\[\frac{ 4 }{ 3 } (3.14)(1860.867)\] now solve this ^^

Nnesha (nnesha):

are you there ?

OpenStudy (anonymous):

5843.12238

Nnesha (nnesha):

1860.867 times 3.14 = ???

OpenStudy (anonymous):

okay

Nnesha (nnesha):

1860.867 times 3.14 = ??? what you get after this ??

Nnesha (nnesha):

5843.12238 i got this i don't how there is 4 ?

Nnesha (nnesha):

which calculator are us using ?

Nnesha (nnesha):

you *

OpenStudy (anonymous):

Thats what i said up there

OpenStudy (anonymous):

look up there

OpenStudy (anonymous):

-.-

Nnesha (nnesha):

5840.4 this one is not same like 5843.12238

OpenStudy (anonymous):

I said 5843.12238 already

OpenStudy (anonymous):

but i thought i was wrong cause you kept asking me the same question

OpenStudy (anonymous):

if you scroll up you can see that said 5843.12238

Nnesha (nnesha):

nvm http://prntscr.com/67e2q6

Nnesha (nnesha):

ohh i thought u said that is final answer :/

Nnesha (nnesha):

multiply that by 4

OpenStudy (anonymous):

23372.48952<------------------ Aswer

OpenStudy (anonymous):

Answer*

Nnesha (nnesha):

divide by 3

OpenStudy (anonymous):

divide what by 6?

Nnesha (nnesha):

by 3

Nnesha (nnesha):

the number u just got divide that by 3

OpenStudy (anonymous):

7790.82984

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Thank you Nnesha

Nnesha (nnesha):

np but listen to me try to play with calculator nobody gonna tell u during final exam that now you have to multiply by this then by this so this question is easy like just for 5 minutes all you have to know the formula and then ur job is to plug correct number into the calculator and then calculator job is to give you answer so becareful when you plug numbers into the calc. best of luck!!! :)

OpenStudy (anonymous):

:) Okay Nnesha i'll try using the calculator less. Let me make ssure i understand with it .. So, i have amy formula and my problem they give me the radius and i plug it in to r, Then when work out (?)^3 i times the answer by 3.14, and THEN i times that answer by 4 THEN i divide my answer with that answer and Done. Correct? xD

Nnesha (nnesha):

i didn't sayz that don't use calculator i said make sure u r plugging number into the calc correctly divide top part with 3